Group Selection Page
HELP
Controls Group
Data Analysis Group
ITDM
WP76
DET
Vacuum
EEE
WP78
WP74
FXE
SCS
SPB
MID
HED
SQS
SXP
Sample Environment
Operation
PSPO
XO
TS
Migrated2Zulip
|
Technical integration and coordination
SASE1 Status and SASE3 Readiness
|
Technical integration and coordination logbook, Contact: A. Violante, G. Wellenreuther |
Not logged in |
 |
|
Message ID: 111
Entry time: 20 Apr 2020, 10:59
|
Author: |
Frederik Wolff-Fabris |
Location: |
General |
Severity: |
Information |
Subject: |
Notes Dispatch Meeting 20.04.2020 |
|
|
General:
- Desy informed a power glitch occurred on Thursday 16th at 04:55AM. TS has not seen any traces of it in the EuXFEL internal GLT.
- From next week the Dispatch meeting will start at 09:30AM.
- An additional Friday's Integration Meeting will happen at 11:30AM. Proper information and invitation will be send soon.
CTRL:
- The notification service for the reduced operation mode is ready for start-up and an e-mail will be send around with instructions.
- Groups/staff will be requested to subscribe to the lists.
- Training will be conducted on next Wednesday at 15:00 and/or next week. One member of each group is required to attend it.
DA:
- Nothing to report. Working busily from home.
EEE:
- No issues to report
VAC:
- Two pumps located in a gas attenuator were OFF due to the power glitch of last Thursday 16/04 at 04:55AM.
XPD
- No issues related to the power glitch were observed.
- XPD has received devices and they were successfully stored in vacuum / clean rooms. Another devices shall arrive this week.
DET:
- Nothing to report.
ITDM:
- Nothing to report.
XRO:
- Nothing to report.
LAS:
- Not present.
SPB:
- Nothing to report.
FXE:
- Not present.
MID:
- Everything is fine.
HED:
- All is fine, all is running.
SCS:
- Not present.
SQS:
- Not present. |